# DynoSAM: Open-Source Smoothing and Mapping Framework for Dynamic SLAM

### v2

@@ VIII-A Dynamic Object Reconstruction
Many reconstruction systems assume the environment to be static and rigid [70, 71], as with many traditional SLAM systems. Leveraging DynoSAM for dynamic object segmentation and motion estimation, we have developed DynORecon [8], a reconstruction system that focuses on incrementally mapping dynamic rigid-body objects. DynORecon employs a submap structure to separate the reconstruction of each dynamic object from the static map, ensuring that object motion does not leave any residual artifacts in the static map.
@@
Figure 10: One of our downstream applications, DynORecon [8], incrementally builds up surface mesh reconstructions of both dynamic objects and static background in the OMD experiment [22]. (a): a visualization of dynamic object meshes and their trajectories (uniquely coloured) as well as the static background (grey), in addition to camera pose and trajectory (green); (b-d): the incrementally constructed mesh of Object $3$.
@@
The accurate motion estimation from DynoSAM facilitates an easy and efficient integration of new object measurements into each existing object reconstruction while maintaining object rigidity. Based on Eq. (10), applying $\prescript{W}{k-1}{\mathbf{H}}_{k}$ as estimated by DynoSAM to all points on the object guarantees a consistent motion from time-step $k-1$ to $k$. Similarly, Eq. (2) and Eq. (4) show that the same motion can consistently transform any reference frame fixed to the object. Similar to the treatment of object pose in Sec. V-B and Sec. V-C, we can define an arbitrary reference frame that is rigidly attached to the object body without any prior knowledge on its pose, and use $\prescript{W}{k-1}{\mathbf{H}}_{k}$ to move this body-fixed frame with respect to the global world frame. This allows us to express all object points in the body-fixed frame, $\prescript{L}{}{\mathbf{m}}$, where they are static and time-invariant with respect to their local frame. This representation allows new observations to be integrated into each object reconstruction while remaining consistent with previous measurements [8].
@@
Fig. 10 presents an example of DynORecon incrementally constructing all $4$ dynamic objects (free-floating cubes) in addition to the static background using DynoSAM’s estimations in the OMD experiment [22]. As shown in Fig. 10 (b-d), accurate motion estimations from the upstream Dynamic SLAM system enables a consistent incremental surface reconstruction as the object undergoes complex movement. Building up a correct representation of dynamic objects, as more of them are observed, provides a more comprehensive understanding of moving obstacles in the scene, and is therefore beneficial to other robotic applications such as planning and navigation.
VIII-B Object Trajectory Prediction
@@
Figure 11: Our second downstream application predicts object trajectories based on the output estimation of DynoSAM. We show inferred predictions compared to the future ground truth trajectory of Object $2$ on KITTI $00$. Three snapshots are shown along the object’s trajectory where (a) illustrates an initial prediction based on only $1$ previous state and (b-c) show predictions based on $4$ prior states.
@@
Trajectory prediction plays a key role in path planning and control in dynamic environments. Safer and more efficient navigation can be achieved by incorporating the anticipated future movements of dynamic objects into navigation algorithms [10, 72, 73, 74].
@@
To test DynoSAM’s usability for a downstream task of trajectory prediction, we integrated our pipeline with Trajectron++ [75], a state-of-the-art trajectory prediction algorithm that uses a graph-structured recurrent model to forecast trajectories. Trajectron++ uses current and historical state information, including position, orientation and velocity for each object in the scene to generate real-time 2D predictions.
To provide this information for training and testing, we utilize the object trajectory data produced by DynoSAM. Object positions are taken directly from DynoSAM’s pose estimates $\prescript{W}{}{\mathbf{L}}_{k}$, while velocities are calculated from the estimated motion $\prescript{W}{k-1}{\mathbf{H}}_{k}$. This processed Dynamic SLAM data is used to train and test Trajectron++ on the KITTI tracking dataset (Sec. VII-B). Our resulting testing dataset included $88$ input instances from KITTI $00$, while use used $406$ training instances collected from the remaining seven KITTI sequences. To illustrate the effectiveness of our methods in a simplified setting, we limit our dataset to only include objects categorized as vehicles. We trained the base Trajectron++ model from scratch within the adaptive prediction framework [76].
@@
Fig. 11 illustrates three snapshots of Object $2$’s (cyclist) predicted trajectory at frames $67$, $85$ and $103$ (a-c respectively). On average, our predictions closely track the ground truth trajectory without exhibiting significant overshooting. These results, while preliminary, demonstrate that estimates from our Dynamic SLAM framework are accurate enough to be successfully used for trajectory predictions tasks. While the achieved predictions could be further improved by incorporating map generated by the DynORecon [8] module into Trajectron++; however, our predictions are already of high accuracy and could be leveraged to improve the motion planning module [9, 74, 77], or as informative motion priors within our Dynamic SLAM framework. Furthermore, the output from DynoSAM is 3D and future work will investigate trajectory forecasting algorithms that generate 3D predictions.
IX Limitations and Future Work
Our proposed method for Dynamic SLAM achieves high accuracy in motion and object trajectory estimation, but the back-end optimization is currently solved in full-batch mode at each step. While the front-end can deliver real-time performance, this paper highlights the vital role that the back-end plays in producing highly accurate and smooth object trajectories. Therefore, our immediate goal is to develop incremental solvers tailored to the specific challenges of the Dynamic SLAM problem.
@@ IX Limitations and Future Work
Currently, our system uses a simplistic approach to modeling smooth object motion changes, relying on a constant motion constraint. As discussed in [17], this is akin to a constant velocity model, i.e. zero acceleration, which produces locally plausible motion but cannot capture motions of accelerating objects as accurately. Moving forward, we will be exploring more advanced motion models, potentially through learning, to better capture the short-term dynamics of each object.
The initial trajectory prediction results shown in Sec. VIII-B are promising; therefore we plan to further integrate these predictions into our system. Additionally, we aim to investigate how understanding the dynamics of each object can provide valuable feedback to enhance the SLAM pipeline and improve visual odometry.
@@
Finally, we plan to leverage advanced transformer models such as SAM2 [78, 79] and FlowFormer [80] to improve scene understanding and explore new directions in Dynamic SLAM. We have designed DynoSAM with an emphasis on modularity and extensibility to facilitate such further research.
X Conclusion
@@
We have introduced DynoSAM, a cutting-edge, open-source framework for Dynamic SLAM that represents a significant advancement in the robotics field. By outlining the key theoretical concepts and formulations underpinning our approach, we provide a robust foundation for tackling the challenges of dynamic environments. DynoSAM offers a well-structured platform for implementing, testing, and evaluating Dynamic SLAM solutions, empowering researchers and practitioners to develop and benchmark innovative methodologies with greater ease and precision. Importantly, our implementation is designed for flexibility, with clearly defined interfaces between modules, facilitating integration with existing and new methods. This contribution paves the way for more reliable and adaptable robotic systems in dynamic and complex settings.
@@ Acknowledgements
We would like to thank Kevin Judd and Jonathan Gammell for providing the results of MVO and their discussions. This research is funded with the support of ARIA Research and the Australian Government via the Department of Industry, Science, and Resources CRC-P program (CRCPXI000007).
@@
As discussed in the work of Zhang et al[15], for any point on a rigid-body object measured in the world frame $\prescript{W}{}{\mathbf{m}}_{k}$, there is the following equation:
@@
where $\prescript{W}{k-1}{\mathbf{H}}_{k}:=\prescript{W}{}{\mathbf{L}}_{k-1}\>\prescript{L_{k-1}}{k-1}{\mathbf{H}}_{k}\>\prescript{W}{}{\mathbf{L}}_{k-1}^{-1}$, an operation referred to as a frame change of a pose transformation by Chirikjian et al[23].
@@
As discussed in [11], we can show that if the body-fixed frame pose change is constant then the absolute reference frame change is constant too:
$\prescript{\prescript{}{}{\mathbf{L}}_{k-1}}{k-1}{\mathbf{H}}_{k}$ $=C=\prescript{\prescript{}{}{\mathbf{L}}_{k}}{k}{\mathbf{H}}_{k+1}$ (43)
$\prescript{W}{k-1}{\mathbf{H}}_{k}$ $=\prescript{W}{}{\mathbf{L}}_{k-1}\>C\>\prescript{W}{}{\mathbf{L}}_{k-1}^{-1}$
$\prescript{W}{k}{\mathbf{H}}_{k+1}$ $=\prescript{W}{}{\mathbf{L}}_{k}\>C\>\prescript{W}{}{\mathbf{L}}_{k}^{-1}$
[…]
$\prescript{W}{k}{\mathbf{H}}_{k+1}$ $=\prescript{W}{}{\mathbf{L}}_{k-1}\>C\>C\>C^{-1}\>\prescript{W}{}{\mathbf{L}}_{k-1}^{-1}$
$=\prescript{W}{}{\mathbf{L}}_{k-1}\>C\>\prescript{W}{}{\mathbf{L}}_{k-1}^{-1}$
@@
This equations shows that when the change in local motion is constant, the corresponding change in absolute motion is also constant. However, when the change in local motion is not constant, the change in absolute motion actually scales with distance the object is from $\{W\}$. This relationship is highly relevant relevant for the object smoothing factor in Eq. (18) which enforces a constant motion constraint in the $\{W\}$ frame. Let the change in local motion be defined as:
${}^{L}\mathbf{C}=\prescript{\prescript{}{}{\mathbf{L}}_{k}}{k}{\mathbf{H}}_{k+1}^{-1}\>\prescript{\prescript{}{}{\mathbf{L}}_{k+1}}{k+1}{\mathbf{H}}_{k+2}\in\mathrm{SE}(3)$ (44)
@@
${}^{W}\mathbf{C}=\prescript{W}{k}{\mathbf{H}}_{k+1}^{-1}\>\prescript{W}{k+1}{\mathbf{H}}_{k+2}\in\mathrm{SE}(3)$ (45)
@@
$=\prescript{W}{}{\mathbf{L}}_{k+1}\>\prescript{W}{}{\mathbf{L}}_{k}^{-1}\>\prescript{W}{}{\mathbf{L}}_{k+1}\>^{L}\mathbf{C}\>\prescript{W}{}{\mathbf{L}}_{k+1}^{-1}$
$=\prescript{W}{k}{\mathbf{H}}_{k+1}\>\prescript{W}{}{\mathbf{L}}_{k+1}\>^{L}\mathbf{C}\>\prescript{W}{}{\mathbf{L}}_{k+1}^{-1}$
Substituting Eq. (46) into Eq. (45) defines the relationship between the change in motion when represented in the local-body and world frames:
${}^{W}\mathbf{C}=\prescript{W}{}{\mathbf{L}}_{k+1}\>^{L}\mathbf{C}\>\prescript{W}{}{\mathbf{L}}_{k+1}^{-1}$ (47)
Consequently, the residual and covariance of the object smoothing factor will increase proportionally with the object’s pose. To mitigate this effect for our smaller-scale experiments, we set the covariance of the smoothing factor to a sufficiently large value.

### v3

@@ VII-A Dynamic Object Reconstruction
Many reconstruction systems assume the environment to be static and rigid [Niessner2013tog, Kahler2015tvcg_infinitam], as with static SLAM systems. Leveraging DynoSAM for object segmentation and motion estimation, we show that the output of DynoSAM can be directly used by a reconstruction system, such as DynORecon [Wang2025icra_dynorecon], to incrementally map dynamic rigid-body objects.
@@
Figure 11: Our downstream application, DynORecon [Wang2025icra_dynorecon], incrementally builds up surface mesh reconstructions of both dynamic objects and static background in the OMD (S4U) experiment [Judd19ral]. (a): dynamic object meshes and their trajectories (uniquely colored) as well as the static background (grey), in addition to camera pose and trajectory (green); (b-d): the incrementally constructed mesh of Object 33.
@@
The accurate motion estimation from DynoSAM facilitates an easy and efficient integration of new object measurements into each existing object reconstruction while maintaining object rigidity. Based on (9), applying $\prescript{W}{k-1}{\mathbf{H}}_{k}$ as estimated by DynoSAM to all points on the object guarantees a consistent motion from time-step k1k-1 to $k$. Equation (2) and (4) further show that the same motion can consistently transform any reference frame fixed to the object. Similar to the treatment of object pose in Sec. IV-B and Sec. IV-C, we can define an arbitrary reference frame that is rigidly attached to the object body without any prior knowledge of its pose, and use $\prescript{W}{k-1}{\mathbf{H}}_{k}$ to move this body-fixed frame with respect to the global world frame. This allows us to express all object points in the body-fixed frame, $\prescript{L}{}{\mathbf{m}}$, where they are static and time-invariant with respect to their local frame. This representation allows new observations to be integrated into each object reconstruction while remaining consistent with previous measurements [Wang2025icra_dynorecon].
@@
Fig. 11 presents an example of DynORecon incrementally constructing all $4$ dynamic objects (free-floating cubes) in addition to the static background using DynoSAM’s estimations in the OMD experiment [Judd19ral]. As shown in Fig. 11 (b-d), accurate motion estimations from the upstream Dynamic SLAM system enables a consistent incremental surface reconstruction as the object undergoes complex movement. Building up a correct representation of dynamic objects, as more of them are observed, provides a more comprehensive understanding of moving obstacles in the scene, and is therefore beneficial to other robotic applications such as planning and navigation.
VII-B Object Trajectory Prediction
@@
Figure 12: Predicted object trajectories based on the output estimation of DynoSAM. We show inferred predictions compared to the future ground truth trajectory of Object 22 on KITTI $00$. Two snapshots are shown along the object’s trajectory where (a) illustrates an initial prediction based on only 11 previous state and (b) shows prediction based on 44 prior states.
@@
Trajectory prediction plays a key role in path planning and control in dynamic environments. Safer and more efficient navigation can be achieved by incorporating the anticipated future movements of dynamic objects into navigation algorithms [Hermann2014, Phillips2011icra_sipp, Finean2021icaps, Finean2023ras].
@@
We therefore integrated our DynoSAM pipeline with Trajectron++ [salzmann2020trajectron++], a state-of-the-art trajectory prediction algorithm that uses a graph-structured recurrent model to forecast trajectories. Trajectron++ uses current and historical state information, including pose and velocity for each object in the scene to generate real-time 2D predictions. Object poses are taken directly from DynoSAM’s pose estimates $\prescript{W}{}{\mathbf{L}}_{k}$, while velocities are calculated from the estimated motion $\prescript{W}{k-1}{\mathbf{H}}_{k}$. This processed DynoSAM output is used to train and test Trajectron++ on the KITTI tracking dataset (Sec. VI-A). We trained the base Trajectron++ model from scratch within the adaptive prediction framework [ivanovic2023AdaptivePrediction].
@@
Fig. 12 presents a snapshot of Object $2$’s (cyclist) predicted trajectory at frames $67$ and $103$ (a-b respectively). On average, our predictions closely track the ground truth trajectory without exhibiting significant overshooting. These results, while preliminary, demonstrate that estimates from our Dynamic SLAM framework are accurate enough for trajectory predictions tasks, and our prediction results are of sufficiently high accuracy to be leveraged for planning modules [Finean2021iros, Finean2023ras, king2022scate], or as informative motion priors within our Dynamic SLAM framework. Furthermore, our future work will investigate trajectory forecasting algorithms that generate 3D predictions fully leveraging the $\mathrm{SE}(3)$ output of DynoSAM.
VIII Limitations and Future Work
Our proposed Dynamic SLAM method achieves high accuracy in both motion and object trajectory estimation, highlighting the crucial role of the back-end in generating precise and smooth object trajectories. While the front-end already operates close to real time, the back-end is currently solved in a full-batch manner and will be improved in future work to support incremental inference.
@@ VIII Limitations and Future Work
Currently, our system uses a simplistic approach that relies on a constant motion constraint to model changes in object motion. As discussed in [judd2024ijrr_mvo], this is akin to a constant velocity model, i.e. zero acceleration, which produces locally plausible motion but cannot capture motions of accelerating objects as accurately. Moving forward, we will exploring more sophisticated motion models, potentially through learning, which better capture an objects the short-term dynamics.
In addition, our system is currently designed to estimate the motion of rigid-bodies. While we have shown examples demonstrating its robustness in semi-rigid cases, a more comprehensive approach to non-rigid cases is highly relevant.
@@
Finally, we plan to leverage advanced transformer models such as SAM2 [ravi2024sam2, yang2024samurai] and FlowFormer [huang2022flowformer] to improve scene understanding and explore new directions in Dynamic SLAM. We have designed DynoSAM with an emphasis on modularity and extensibility to facilitate such further research.
IX Conclusion
@@
We have introduced DynoSAM, a cutting-edge, open-source framework for Dynamic SLAM that represents a significant advancement in the robotics field. By outlining the key theoretical concepts and formulations underpinning our approach, we provide a robust foundation for tackling the challenges of estimation in dynamic environments. DynoSAM offers a well-structured platform for implementing, testing, and evaluating Dynamic SLAM solutions, empowering researchers to develop and benchmark innovative methodologies with greater ease and precision. Importantly, our implementation is designed for flexibility, with clearly define interfaces between modules, facilitating integration with existing and new methods. This contribution paves the way for more reliable and adaptable robotic systems in dynamic and complex settings.
@@ Acknowledgements
We would like to thank Kevin Judd and Jonathan Gammell for providing the results of MVO and their discussions. We additionally thank Tim Bailey for pointing out the scaling behaviour when evaluating world frame motion error.
@@
As discussed in the work of Zhang et al[zhang2020vdoslam], for any point on a rigid-body object measured in the world frame $\prescript{W}{}{\mathbf{m}}_{k}$, there is the following equation:
@@
where $\prescript{W}{k-1}{\mathbf{H}}_{k}:=\prescript{W}{}{\mathbf{L}}_{k-1}\>\prescript{L_{k-1}}{k-1}{\mathbf{H}}_{k}\>\prescript{W}{}{\mathbf{L}}_{k-1}^{-1}$, an operation referred to as a frame change of a pose transformation by Chirikjian et al[Chirikjian17idetc].
@@
As discussed by Henein et al[Henein20icra], we can show that if the body-fixed frame pose change is constant then the absolute reference frame change is constant too:
$\prescript{\prescript{}{}{\mathbf{L}}_{k-1}}{k-1}{\mathbf{H}}_{k}$ $=\mathbf{C}=\prescript{\prescript{}{}{\mathbf{L}}_{k}}{k}{\mathbf{H}}_{k+1}$ (38)
$\prescript{W}{k-1}{\mathbf{H}}_{k}$ $=\prescript{W}{}{\mathbf{L}}_{k-1}\>\mathbf{C}\>\prescript{W}{}{\mathbf{L}}_{k-1}^{-1}$
$\prescript{W}{k}{\mathbf{H}}_{k+1}$ $=\prescript{W}{}{\mathbf{L}}_{k}\>\mathbf{C}\>\prescript{W}{}{\mathbf{L}}_{k}^{-1}$
$\text{given that,}\quad\prescript{W}{}{\mathbf{L}}_{k}$ $=\prescript{W}{}{\mathbf{L}}_{k-1}\>\mathbf{C}$
$\prescript{W}{k}{\mathbf{H}}_{k+1}$ $=\prescript{W}{}{\mathbf{L}}_{k-1}\>\mathbf{C}\>\mathbf{C}\>\mathbf{C}^{-1}\>\prescript{W}{}{\mathbf{L}}_{k-1}^{-1}$
$=\prescript{W}{}{\mathbf{L}}_{k-1}\>\mathbf{C}\>\prescript{W}{}{\mathbf{L}}_{k-1}^{-1}$
@@
However, when the change in local motion is not constant, the change in absolute motion scales with distance the object is from $\{W\}$. This relationship is highly relevant for the object smoothing factor in (17) which enforces a constant motion constraint in the $\{W\}$ frame. Let the change in local motion be defined as:
$^{L}\mathbf{\Delta}=\prescript{\prescript{}{}{\mathbf{L}}_{k}}{k}{\mathbf{H}}_{k+1}^{-1}\>\prescript{\prescript{}{}{\mathbf{L}}_{k+1}}{k+1}{\mathbf{H}}_{k+2}\in\mathrm{SE}(3)$ (39)
@@
$^{W}\mathbf{\Delta}=\prescript{W}{k}{\mathbf{H}}_{k+1}^{-1}\>\prescript{W}{k+1}{\mathbf{H}}_{k+2}\in\mathrm{SE}(3)$ (40)
@@
$=\prescript{W}{}{\mathbf{L}}_{k+1}\>\prescript{W}{}{\mathbf{L}}_{k}^{-1}\>\prescript{W}{}{\mathbf{L}}_{k+1}\>^{L}\mathbf{\Delta}\>\prescript{W}{}{\mathbf{L}}_{k+1}^{-1}$
$=\prescript{W}{k}{\mathbf{H}}_{k+1}\>\prescript{W}{}{\mathbf{L}}_{k+1}\>^{L}\mathbf{\Delta}\>\prescript{W}{}{\mathbf{L}}_{k+1}^{-1}$
Substituting (41) into (40) defines the relationship between the change in motion when represented in the local-body and world frames:
$^{W}\mathbf{\Delta}=\prescript{W}{}{\mathbf{L}}_{k+1}\>^{L}\mathbf{\Delta}\>\prescript{W}{}{\mathbf{L}}_{k+1}^{-1}$ (42)
Consequently, the residual and covariance of the object smoothing factor will increase proportionally with the object’s pose. To mitigate this effect for our smaller-scale experiments, we set the covariance of the smoothing factor to a sufficiently large value. The same ‘scaling’ behavior as represented by (42) is also present when evaluating motion errors using $\prescript{W}{}{\mathbf{H}}$ instead of $\prescript{L}{}{\mathbf{H}}$, as a small discrepancies in the local motion $\prescript{L}{}{\mathbf{H}}$ can lead to a large error in $\prescript{W}{}{\mathbf{H}}$ due to (42).
[Uncaptioned image] Jesse Morris is a 3rd year PhD student at the Australian Centre For Robotics (ACFR), supervised by Dr Viorela Ila. His thesis focuses on developing estimation frameworks for Dynamic SLAM. He received his bachelors degree in Mechatronic Engineering and Computer Science from the University of Sydney in 2022.
[Uncaptioned image] Yiduo Wang is a postdoctoral researcher in the ACFR, working with Dr Viorela Ila. He focuses on SLAM, reconstruction and navigation in dynamic environments. He holds a DPhil Engineering Science degree awarded by University of Oxford, where his research focused on large-scale reconstruction. He also holds an MRes Robotics degree awarded by UCL on combining SLAM with semantic segmentation for dynamic environments.
[Uncaptioned image] Mikolaj Kliniewski is a 2nd year PhD student at the Australian Centre For Robotics, supervised by Dr Viorela Ila. His thesis focuses on joint planning and estimation in dynamic environments. He received his Bachelor Honours degree in Computer Science from the University of Liverpool in 2023.
[Uncaptioned image] Viorela Ila received a Ph.D. in Information Technologies from the Universitat de Girona, Spain. She worked at the Institut de Robótica i Informàtica Industrial, Barcelona, and was awarded a MICINN/FULBRIGHT post-doctoral fellowship in 2009, joining Georgia Tech, USA. She joined the ROSACE project in LAAS-CNRS, France, in 2010, and was a Research Scientist at Brno University of Technology, Czech Republic (2012–2014) and a Research Fellow at the Australian National University (2015–2018). Currently a Senior Lecturer at the University of Sydney, her research focuses on robot vision, SLAM and 3D reconstruction, leveraging graphical models, optimization methods and information theory.